Grade 12Electrostatics

when an electron moves from low potential to high potential ;is its potential energy increase and why?

ankit
12 Years agoGrade 12

1 Answer

dfdffdfddffd
ApprovedApproved Tutor Answer12 Years ago
As a rule every thing in this physical world tries to minimize its "energy" or become less excited. An electron with negative change lying at a distance (x) from a positive plate (+Q) entails PE = - k Q e/x say this value ( = - 10 units) .... suppose as a contradiction electron RUNS away from plate, then (x) will increase, say reaches (2x) .... PE = - 10/2 = - 5 units ........... [- 5 > - 10 ] ......... so in this contradiction electron increases its "energy" ..... this physics denies ........ so electron has the sole option to run towards the plate. 

That is why 2 negative charges REPEL and opposite charges attract [Try yourself bu using energy units like above] .................. hope my answer will thrill you?